OpenStudy (australopithecus):

Why are the following integrals divergent? \[\int\limits_{1}^{\infty}(x+1)dx/(x^{2} +2x)\] \[\int\limits_{2pi}^{\infty}\sin(x)dx\]

OpenStudy (anonymous):

sin(x) is divergent because it does not have a limit, it will be between [-1,1] and the sum will never converge.

OpenStudy (anonymous):

The other one I suspect it's because the antiderivative will be logs. When t -> infinity, the log will diverge and the integral will diverge.

OpenStudy (australopithecus):

x/(x+2) is divergent thus the entire integral is divergent

OpenStudy (australopithecus):

for the first one

OpenStudy (anonymous):

This is what I mean: \[lim_{t \rightarrow \infty} \int_{1}^{t} \frac{x+1}{x^2 + 2x} dx = lim_{t \rightarrow \infty} (\ln{t} + \ln{t+2})\]Can be thought that way too.

OpenStudy (anonymous):

Should be ln(t+2)

OpenStudy (australopithecus):

you can split it